# Optuna vs Code Llama: Are LLMs a New Paradigm for Hyperparameter Tuning?

### v1

@@ Abstract
Optimal hyperparameter selection is critical for maximizing neural network performance, especially as models grow in complexity. This work investigates the viability of using large language models (LLMs) for hyperparameter optimization by employing a fine-tuned version of Code Llama. Through parameter-efficient fine-tuning using LoRA, we adapt the LLM to generate accurate and efficient hyperparameter recommendations tailored to diverse neural network architectures. Unlike traditional methods such as Optuna, which rely on exhaustive trials, the proposed approach achieves competitive or superior results in terms of Root Mean Square Error (RMSE) while significantly reducing computational overhead. Our approach highlights that LLM-based optimization not only matches state-of-the-art methods like Tree-structured Parzen Estimators but also accelerates the tuning process. This positions LLMs as a promising alternative to conventional optimization techniques, particularly for rapid experimentation. Furthermore, the ability to generate hyperparameters in a single inference step makes this method particularly well-suited for resource-constrained environments such as edge devices and mobile applications, where computational efficiency is paramount. The results confirm that LLMs, beyond their efficiency, offer substantial time savings and comparable stability, underscoring their value in advancing machine learning workflows. All generated hyperparameters are included in the LEMUR Neural Network (NN) Dataset, which is publicly available on GitHub and serves as an open-source benchmark for hyperparameter optimization research.
@@ 1 Introduction
It is a well-established fact that the choice of training hyperparameters significantly influences the learning efficiency and accuracy of machine learning algorithms. Several hyperparameter tuning approaches have already been proposed, emphasizing their importance in the learning process. It has been demonstrated that a poor choice of these parameters can lead to suboptimal learning processes and negatively impact the final results, whereas well-informed choices yield significantly better outcomes with greater efficiency using the same model structure.
LLMs have demonstrated the ability to capture complex relationships across various applications, as discussed in Section 1.1. Fine-tuning techniques enable LLMs to acquire new information and solve previously unseen tasks. In this work, we examine state-of-the-art fine-tuning techniques to enhance the capability of LLMs in predicting optimal hyperparameters for neural network models, proposing this as a novel approach to hyperparameter optimization.
This study focuses on training hyperparameters of well-known computer vision models, specifically for the task of image classification.
@@ 1.1 Related work
Early model-free approaches to the to the hyperparameter optimization task (HPO), such as grid and random search, perform well in low-dimensional hyperparameter spaces. However, as the number of hyperparameters increases, the computational cost of applying these methods also rises. For further insights into HPO, we refer readers to [16, 33, 3, 5, 4, 8].
@@
More advanced techniques for HPO include evolutionary strategies (ES) and bayesian optimization (BO). Although ES outperforms earlier methods in more complex search spaces, their application can still be computationally expensive [8, 9].
@@
BO is another HPO approach that has proven to be efficient in high-dimensional hyperparameter spaces. It surpasses ES in terms of computational efficiency. BO works by constructing a surrogate model, updating its posterior distribution based on observed data, and selecting the next candidate using an acquisition function. The efficiency of BO is closely tied to the choice of surrogate model, with the tree-structured Parzen estimator (TPE) being the focus of this study [6]. TPE models the objective function by creating probability distributions for both favorable and less favorable hyperparameter values. As optimization progresses, these distributions are iteratively refined, guiding the search toward regions with a higher probability of yielding optimal outcomes. For a more detailed discussion of BO, we refer to [37, 16, 38, 7, 14, 23].
@@
Applying this framework, Optuna [1] is an adaptable hyperparameter optimization tool that supports various surrogate models within the BO approach. For this study, we compare Optuna’s TPE-based approach, due to its effectiveness in navigating complex hyperparameter spaces, with the Python version of Code Llama (7B and 13B parameter versions) [36].
In this study, we compare Optuna’s TPE-based approach — selected for its effectiveness in navigating complex hyperparameter spaces — with the Python version of Code Llama (7B-parameter model) [36].
@@
The Transformer architecture [40, 13], which led to the development of LLMs, has revolutionized a wide range of fields, such as natural language processing [10, 44, 21], code generation [35, 32, 43, 15, 26], computer vision [39, 41, 2], complex reasoning [20, 45, 24, 29, 18, 46, 42] and many others [34, 47, 22].
@@
Extensive research has been conducted to assess whether LLMs can offer superior results compared to established HPO methods. [48] demonstrated that LLMs could achieve performance comparable to or better than that of conventional methods when applied to models such as logistic regression, support vector machines (SVMs), random forests, and neural networks. Their study compared various HPO algorithms, including BO with Gaussian process and random forest surrogates, to random search baseline. [11] analyzed use of LLMs for tuning the step-size of the (1+1)-ES algorithm, demonstrating that LLM-driven strategies can effectively compete with established methods. [27] introduced AgentHPO, a hyperparameter optimization tool that works iteratively and frequently outperforms human-derived solutions while delivering both high performance and interpretability. [28] proposed LLAMBO, a method that integrates LLMs with traditional BO and has been proven to be an effective zero-shot warm-starting technique by leveraging LLMs’ domain knowledge. Lastly, in a recent study, [30] presented sequential large language model-based optimization (SLLMBO), which utilizes LLMs for HPO, showcasing the possibility of outperforming traditional BO methods.
The distinguishing feature of this study, compared to recent works, is the fine-tuning process applied to the chosen LLM. To distinguish our work from that of [30], we emphasize that the models used in our study are locally deployable. We recognize that there is potential for improvement upon our work due to resource limitations. However, to the best of our knowledge, the findings of our research have not been presented before.
@@ 4.3 Model Performance Analysis
Models from train dataset. The comparison between models included in fine-tuning along with their 5-epoch versions (AlexNet, RNN, and ConvNeXt) and models whose 5-epoch versions were excluded from fine-tuning (ResNet, RegNet, MobileNetV2) highlights the impact of fine-tuning on RMSE values. As shown in Figure 3(a), models that underwent fine-tuning demonstrate consistently lower RMSE values, with significantly reduced confidence intervals, reflecting improved stability and robustness in hyperparameter recommendations. ConvNeXt, in particular, exhibits a notable decrease in RMSE, confirming the effectiveness of fine-tuning. Models whose 5-epoch versions were not used in fine-tuning (Figure 3(b)) generally exhibit higher RMSE values and broader confidence intervals, indicating increased variability and reduced optimization quality. Notably, ResNet shows a slight increase in RMSE compared to the Optuna baseline, suggesting that fine-tuning did not yield improvements for this particular architecture. However, other models in this group, such as MobileNetV2, still benefit from fine-tuning, with confidence intervals narrowing and RMSE values showing moderate reductions. These results highlight the nuanced impact of fine-tuning, where some architectures see substantial gains while others may require additional adjustments to maximize improvements.
@@ 4.4 Prediction Dynamics Across Epochs
In this experiment, we analyze hyperparameter recommendations for a diverse set of models. While the full evaluation was conducted across all models in our study, we present detailed results for four representative architectures: MobileNetV2 and DenseNet, which were part of the training set (participated in Code Llama fine-tuning with data from 1 and 2 epochs), and MaxVit and VisionTransformer, which belong to the test set (not involved in Code Llama fine-tuning). This selection provides a balanced comparison between models seen and unseen during fine-tuning. The experiment illustrates the accuracy of one-shot predictions across multiple epochs (1, 2, 5, 10, 15, and 20) obtained using fine-tuned Code Llama compared to the best and average accuracy values achieved using Optuna trials. The Figure 4 shows that the training models (MobileNetV2 and DenseNet) demonstrate a significant increase in the accuracy of one-shot predictions with the increasing number of epochs, exceeding the average values of Optuna and approaching its best results. In the case of the test models (MaxVit and VisionTransformer), one-shot predictions also show impressive results. This is especially noticeable for the MaxVit model, where the one-shot accuracy reaches around 0.7, while the best accuracy of Optuna is only 0.1. These results highlight that Code Llama effectively tunes hyperparameters and provides high prediction accuracy, even for models that have not been trained, indicating a strong generalization ability of the method. While exhaustive search can theoretically yield optimal hyperparameter values, the associated time costs are a considerable constraint. In real-world model training, exhaustive search becomes highly impractical, underscoring the value of efficient hyperparameter tuning methods like Code Llama.
@@ 4.5 Training Trends and Performance Comparison
Optuna, as seen from the green dots in the graphs, shows a significant spread in accuracy values, especially for a model like ConvNeXt, indicating high variability in results. At the same time, fine-tuning shows more stable results, confirming its effectiveness in achieving consistent accuracy. The greatest effect of fine-tuning is seen for ResNet and DenseNet, where the blue line is noticeably higher than the purple and green ones, indicating significant improvements. For some models, like GoogLeNet, the difference between the best values after fine-tuning and the one-shot predictions is less pronounced, indicating limited potential for improvement for individual architectures, but even here fine-tuning maintains a small advantage in accuracy.
@@ 5 Conclusion
Moreover, by reducing the computational cost of hyperparameter tuning, our approach aligns with the objectives of efficient and resource-aware AI, making it particularly relevant for deployment on edge devices. The ability to optimize models with minimal overhead supports the broader goal of enabling lightweight AI applications in constrained environments such as mobile phones and virtual headsets.
@@ 6.1 Models from train dataset
AlexNet 1 0.2397 0.2154 0.0243 0.4359 -0.1962
AlexNet 2 0.2550 0.4253 -0.1704 0.5829 -0.3279
ConvNeXt 1 0.2227 0.2563 -0.0336 0.2516 -0.0289
ConvNeXt 2 0.2780 0.3188 -0.0408 0.3183 -0.0402
DenseNet 1 0.5252 0.5644 -0.0392 0.4705 0.0546
DenseNet 2 0.6656 $0$0.0344 0.6401 0.0255
EfficientNet 1 0.5309 0.3986 0.1323 0.3711 0.1598
EfficientNet 2 0.6752 0.4953 0.1799 0.4768 0.1984
GoogLeNet 1 0.5047 0.4812 0.0235 0.5351 -0.0305
GoogLeNet 2 0.6301 0.6481 -0.0180 0.6980 -0.0679
MNASNet 1 0.5557 0.4775 0.0782 0.5018 0.0538
MNASNet 2 0.6655 0.5823 0.0832 0.5504 0.1151
MobileNetV2 1 0.5527 0.5938 -0.0411 0.5843 -0.0315
MobileNetV2 2 0.6665 0.7152 -0.0487 0.6079 0.0586
MobileNetV3 1 0.5446 0.3101 0.2345 0.2190 0.3256
MobileNetV3 2 0.6788 0.3668 0.3121 0.3653 0.3136
RegNet 1 0.4448 0.4679 -0.0231 0.3934 0.0514
RegNet 2 0.5790 0.6144 -0.0354 0.4768 0.1022
ResNet 1 0.5396 0.5089 0.0307 0.4595 0.0801
ResNet 2 0.6766 0.6117 0.0649 0.4939 0.1827
ShuffleNet 1 0.5412 0.5838 -0.0426 0.5108 0.0304
ShuffleNet 2 0.6617 0.6877 -0.0260 0.5889 0.0727
SqueezeNet 1 0.1669 0.2886 -0.1217 0.3136 -0.1468
SqueezeNet 2 0.1931 0.3567 -0.1637 0.4198 -0.2268
SwinTransformer 1 0.2861 0.3529 -0.0668 0.3287 -0.0426
SwinTransformer 2 0.3830 0.4426 -0.0595 0.3832 -0.0002
VGG 1 0.2616 0.4503 -0.1888 0.4670 -0.2054
VGG 2 0.3323 0.6279 -0.2957 0.6268 -0.2946
@@
Table 2: Comparison of RMSE differences for fine-tuning cycles 1 and 2 for each computer vision model and epoch from the training dataset. The table highlights the RMSE values obtained using Optuna as a baseline (RMSE Optuna) and compares them with the RMSE values achieved during the first and second fine-tuning cycles of Code Llama (RMSE FT Cycle 1 and RMSE FT 2, respectively). The Difference Cycle 1 and Difference Cycle 2 columns indicate the changes in RMSE relative to Optuna for the first and second fine-tuning cycles, respectively. Positive values in the difference columns represent improvements (i.e., reductions in RMSE), while negative values indicate an increase in RMSE.
Model RMSE Optuna RMSE FT 1 Difference
RNN 0.2485 0.2500 -0.0015
LSTM 0.0023 0.0017 0.0006
Llama 0.2144 0.2251 -0.0107
@@
Table 3: Comparison of RMSE values for text generation models from the training dataset during fine-tuning cycle 1. The table highlights the differences in RMSE between the baseline results obtained using Optuna and the fine-tuned results (FT 1). Positive differences indicate an improvement in RMSE after fine-tuning, while negative differences suggest a potential decline in performance.
Computer vision models. The analysis of the results from fine-tuning Code Llama across cycle 1 and cycle 2, as shown in Table 2, explores the impact of iterative fine-tuning on a diverse set of computer vision models compared to baseline performance achieved using Optuna.
In the first fine-tuning cycle, several models demonstrate strong improvements in RMSE compared to the Optuna baseline. For instance, EfficientNet exhibits a substantial reduction in RMSE across both epochs (1st epoch: +0.1323, 2nd epoch: +0.1799), while MobileNetV3 achieves remarkable gains (1st epoch: +0.2345, 2nd epoch: +0.3121). These results underline the effectiveness of fine-tuning for models with a high degree of adaptability to this process. Similarly, ResNet benefits from fine-tuning, showing consistent positive differences in both epochs during the first cycle (e.g., 2nd epoch: +0.0649). Other architectures, such as DenseNet, exhibit stable performance with slight reductions in RMSE compared to Optuna.
The second fine-tuning cycle further enhances the performance for certain models, particularly those that had already demonstrated improvements in the first cycle. For example, MobileNetV3 continues to show notable positive differences in RMSE during both epochs (1st epoch: +0.3256, 2nd epoch: +0.3136), indicating that additional iterations of fine-tuning can extract further gains for architectures that are highly receptive to optimization. Similarly, EfficientNet shows continued improvements in both epochs (1st epoch: +0.1598, 2nd epoch: +0.1984), reflecting the iterative benefits of fine-tuning. For some models, such as DenseNet, the second cycle yields an improvement over the first (e.g., 1st epoch: +0.0546), suggesting that additional fine-tuning can mitigate initial reductions in RMSE.
However, not all models benefit equally from iterative fine-tuning. For instance, SqueezeNet demonstrates an increasing RMSE difference in the second epoch (-0.2268), suggesting limited adaptability to the fine-tuning process in this context. Similarly, GoogLeNet shows minor performance drops during the second cycle (2nd epoch: -0.0679), highlighting the variability in fine-tuning effectiveness across architectures. This underscores the need to carefully evaluate the trade-offs and potential limitations when applying fine-tuning to certain model types.
@@
The comparison between fine-tuning cycles and the Optuna baseline reveals that iterative fine-tuning effectively reduces RMSE for the majority of models and epochs, particularly for those architectures that show high adaptability, such as EfficientNet, MobileNetV3, and ResNet. These findings provide evidence that fine-tuning is a robust optimization strategy capable of surpassing or complementing traditional hyperparameter search methods like Optuna for many computer vision tasks. Furthermore, the variability observed across models and epochs emphasizes the importance of tailoring fine-tuning strategies to the specific characteristics of individual architectures to maximize performance.
@@
Text generation models. The analysis of the RMSE differences for fine-tuning cycle 1, as shown in Table 3, provides insights into the performance impact of fine-tuning on text generation models.
For the RNN model, the RMSE after fine-tuning showed a minimal increase of 0.0015 compared to the baseline obtained using Optuna. This suggests that fine-tuning had a negligible effect on the performance of the model, with results remaining stable. The slight increase could indicate a marginal loss in generalization, but it does not significantly affect the model’s reliability.
The LSTM model, on the other hand, exhibited a slight improvement in RMSE, with a reduction of 0.0006 after fine-tuning. While this improvement is minor, it demonstrates that fine-tuning was able to optimize the model’s parameters effectively, albeit with a minimal effect. This stability and slight enhancement underscore the robustness of LSTM architectures to fine-tuning adjustments.
In contrast, the Llama model experienced the largest increase in RMSE, with a difference of -0.0107 compared to the baseline. This indicates that fine-tuning in this case resulted in a decrease in performance. The deterioration may suggest challenges in adapting the Llama architecture to the fine-tuning process or potential overfitting to the specific fine-tuning dataset. These results emphasize the need for careful calibration of fine-tuning strategies for more complex architectures like Llama.

### v2

@@ Abstract
Optimal hyperparameter selection is critical for maximizing neural network performance, especially as models grow in complexity. This work investigates the viability of leveraging large language models (LLMs) for hyperparameter optimization by fine-tuning a parameter-efficient version of Code Llama using LoRA. The adapted LLM is capable of generating accurate and efficient hyperparameter recommendations tailored to diverse neural network architectures. Unlike traditional approaches such as Optuna, which rely on computationally intensive trial-and-error procedures, our method achieves competitive or superior results in terms of Root Mean Square Error (RMSE) while significantly reducing computational overhead. Our findings demonstrate that LLM-based optimization not only matches the performance of state-of-the-art techniques like Tree-structured Parzen Estimators (TPE) but also substantially accelerates the tuning process. This positions LLMs as a promising alternative for rapid experimentation, particularly in resource-constrained environments such as edge devices and mobile platforms, where computational efficiency is essential. In addition to improved efficiency, the method offers time savings and consistent performance across various tasks, highlighting its robustness and generalizability. All generated hyperparameters are included in the LEMUR Neural Network (NN) Dataset, which is publicly available and serves as an open-source benchmark for hyperparameter optimization research.
@@ 1 Introduction
The choice of training hyperparameters plays a crucial role in determining the learning efficiency and performance of machine learning models. Poorly selected hyperparameters can lead to suboptimal learning and degraded accuracy, while well-chosen configurations can significantly boost performance using the same model architecture.
Large Language Models (LLMs) have recently demonstrated the ability to capture complex patterns across diverse tasks. With fine-tuning techniques, LLMs can acquire new capabilities and solve previously unseen problems. In this work, we explore fine-tuning LLMs for hyperparameter optimization (HPO), positioning them as a novel alternative to traditional methods such as Bayesian Optimization (BO) and Evolutionary Strategies (ES).
Our approach leverages a fine-tuned version of Code Llama to recommend hyperparameters for deep learning models in a one-shot manner. We focus on training hyperparameters for well-known computer vision architectures in the context of image classification, aiming to evaluate whether LLMs can match or surpass the performance of traditional optimization frameworks such as Optuna.
@@ 1.1 Related work
Early model-free approaches to the hyperparameter optimization (HPO) task, such as grid and random search, perform well in low-dimensional hyperparameter spaces. However, as the number of hyperparameters increases, the computational cost of applying these methods also rises. For further insights into HPO, we refer readers to [16, 33, 3, 5, 4, 8].
@@
More advanced techniques for HPO include evolutionary strategies (ES) and Bayesian optimization (BO). Although ES outperforms earlier methods in more complex search spaces, its application can still be computationally expensive [8, 9].
@@
BO has proven to be an efficient alternative, especially in high-dimensional hyperparameter spaces, and typically surpasses ES in terms of computational efficiency. It operates by constructing a surrogate model, updating its posterior distribution based on observed data, and selecting the next candidate using an acquisition function. The effectiveness of BO strongly depends on the surrogate model used, with the tree-structured Parzen Estimator (TPE) being the focus of this study [6]. TPE models the objective function by creating probability distributions for both favorable and unfavorable hyperparameter values. These distributions are iteratively refined during optimization, guiding the search toward regions with a higher likelihood of success. For a more detailed discussion of BO, we refer to [37, 16, 38, 7, 14, 23].
@@
Applying this framework, Optuna [1] is a flexible hyperparameter optimization tool that supports various surrogate models within the BO paradigm. For this study, we compare Optuna’s TPE-based approach—due to its effectiveness in navigating complex hyperparameter spaces—with the Python version of Code Llama (7B and 13B parameter versions) [36].
@@
The Transformer architecture [40, 13], which underlies most LLMs, has revolutionized a wide range of fields, including natural language processing [10, 44, 21], code generation [35, 32, 43, 15, 26], computer vision [39, 41, 2], and complex reasoning tasks [20, 45, 24, 29, 18, 46, 42], among many others [34, 47, 22].
@@
Extensive research has been conducted to evaluate whether LLMs can outperform established HPO methods. For instance, [48] demonstrated that LLMs can achieve performance comparable to or better than conventional methods when applied to models such as logistic regression, support vector machines (SVMs), random forests, and neural networks. Their study compared various HPO algorithms, including BO with Gaussian process and random forest surrogates, to a random search baseline. [11] investigated using LLMs to tune the step-size of the (1+1)-ES algorithm, showing that LLM-driven strategies can effectively compete with traditional methods. [27] introduced AgentHPO, an iterative HPO tool powered by LLMs, which frequently outperforms human-crafted solutions while maintaining interpretability. [28] proposed LLAMBO, a method that integrates LLMs with BO, demonstrating effective zero-shot warm-starting by leveraging LLMs’ domain knowledge. Lastly, [30] presented Sequential Large Language Model-Based Optimization (SLLMBO), showcasing how LLMs can outperform traditional BO methods in several scenarios.
The distinguishing feature of our study is the application of fine-tuning to the LLM for the HPO task. Unlike previous work such as [30], we focus on locally deployable models. While we acknowledge that our setup has resource limitations, to the best of our knowledge, the findings and methodology presented here are novel and have not been previously explored.
@@ 4.3 Model Performance Analysis
Models from train dataset. The comparison between models included in fine-tuning along with their 5-epoch versions (AlexNet, RNN, and ConvNeXt) and models whose 5-epoch versions were excluded from fine-tuning (ResNet, RegNet, MobileNetV2) highlights the impact of fine-tuning on RMSE values. As shown in Figure 3(a), models that underwent fine-tuning demonstrate consistently lower RMSE values, with significantly reduced confidence intervals, reflecting improved stability and robustness in hyperparameter recommendations. ConvNeXt, in particular, exhibits a notable decrease in RMSE, confirming the effectiveness of fine-tuning. Models whose 5-epoch versions were not used in fine-tuning (Figure 3(b)) exhibit higher RMSE values in the LLM’s predictions compared to Optuna, indicating room for improvement in generalization. Nonetheless, the fine-tuned model maintains reasonable stability, as reflected in confidence intervals that, while broader, still reflect consistent behavior across trials.
@@ 4.4 Prediction Dynamics Across Epochs
In this experiment, we analyze hyperparameter recommendations for a diverse set of models. While the full evaluation was conducted across all models in our study, we present detailed results for four representative architectures: MobileNetV2 and DenseNet, which were part of the training set (participated in Code Llama fine-tuning with data from 1 and 2 epochs), and MaxVit and VisionTransformer, which belong to the test set (not involved in Code Llama fine-tuning). This selection provides a balanced comparison between models seen and unseen during fine-tuning.
The experiment illustrates the accuracy of one-shot predictions across multiple epochs (1, 2, 5, 10, 15, and 20), as obtained using the fine-tuned Code Llama, and compares them to the best and average accuracy values achieved by Optuna. As shown in Figure 4, training models (MobileNetV2 and DenseNet) exhibit a consistent increase in prediction accuracy as the number of epochs increases. Their performance surpasses the average Optuna accuracy and closely approaches the best Optuna results, confirming the effectiveness of fine-tuning on seen models.
For test models (MaxVit and VisionTransformer), Code Llama also demonstrates strong generalization. Notably, MaxVit achieves one-shot accuracy close to 0.7, significantly outperforming both the average and best Optuna results, which remain at 0.1. VisionTransformer also steadily improves with more training epochs, showing superior performance relative to Optuna’s average and best scores. Detailed results for additional models can be found in the Supplementary Material (Figures 67). These outcomes highlight Code Llama’s ability to generate competitive hyperparameter configurations even for unseen architectures.
While exhaustive search can theoretically yield optimal hyperparameter values, the associated time costs are a considerable constraint. In real-world model training, exhaustive search becomes highly impractical, underscoring the value of efficient hyperparameter tuning methods like Code Llama.
@@ 4.5 Training Trends and Performance Comparison
Optuna, as seen from the green dots in the graphs, shows a significant spread in accuracy values, especially for a model like ConvNeXt and MobileNetV3, indicating high variability in results. At the same time, fine-tuning shows more stable results, confirming its effectiveness in achieving consistent accuracy. The greatest effect of fine-tuning is seen for DenseNet and MobileNetV3, where the blue line is noticeably higher than the purple and green ones, indicating significant improvements. For some models, like GoogLeNet, the difference between the best values after fine-tuning and the one-shot predictions is less pronounced, indicating limited potential for improvement for individual architectures, but even here fine-tuning maintains a small advantage in accuracy. Detailed results for additional models can be found in the Supplementary Material (Figures 8 - 10).
@@ 6.1 Models from train dataset
AlexNet 1 0.8258 0.8421 -0.0163 0.5779 0.2479
AlexNet 2 0.8236 0.6912 0.1324 0.4221 0.4015
ConvNeXt 1 0.7924 0.7460 0.0464 0.7493 0.0431
ConvNeXt 2 0.7426 0.6861 0.0565 0.6863 0.0563
DenseNet 1 0.4930 0.4526 0.0404 0.5899 -0.0969
DenseNet 2 0.3525 0.3325 0.0201 0.4307 -0.0782
EfficientNet 1 0.4835 0.6250 -0.1415 0.6498 -0.1663
EfficientNet 2 0.3370 0.5486 -0.2116 0.5563 -0.2193
GoogLeNet 1 0.5158 0.5865 -0.0707 0.5093 0.0065
GoogLeNet 2 0.3956 0.3857 0.0099 0.3044 0.0912
MNASNet 1 0.4613 0.5309 -0.0696 0.5132 -0.0519
MNASNet 2 0.3682 0.4308 -0.0626 0.4620 -0.0938
MobileNetV2 1 0.4654 0.4155 0.0499 0.4316 0.0338
MobileNetV2 2 0.3617 0.2854 0.0763 0.4279 -0.0662
MobileNetV3 1 0.4852 0.7328 -0.2476 0.8186 -0.3334
MobileNetV3 2 0.3368 0.6797 -0.3429 0.6567 -0.3199
RegNet 1 0.5677 0.5392 0.0285 0.6254 -0.0577
RegNet 2 0.4368 0.3865 0.0503 0.5484 -0.1116
ResNet 1 0.4692 0.5153 -0.0461 0.5847 -0.1155
ResNet 2 0.3371 0.4325 -0.0954 0.5601 -0.2230
ShuffleNet 1 0.4832 0.4282 0.0550 0.5222 -0.0390
ShuffleNet 2 0.3471 0.3236 0.0235 0.4638 -0.1167
SqueezeNet 1 0.8597 0.7200 0.1397 0.6888 0.1709
SqueezeNet 2 0.8532 0.6676 0.1856 0.5829 0.2703
SwinTransformer 1 0.7500 0.6527 0.0973 0.6746 0.0754
SwinTransformer 2 0.6726 0.5676 0.1050 0.6237 0.0489
VGG 1 0.8192 0.5898 0.2295 0.5364 0.2829
VGG 2 0.7961 0.3741 0.4220 0.3743 0.4218
@@
Table 2: Comparison of RMSE differences for fine-tuning cycles 1 and 2 for each computer vision model and epoch from the training dataset. The table highlights the RMSE values obtained using Optuna as a baseline (RMSE Optuna) and compares them with the RMSE values achieved during the first and second fine-tuning cycles of Code Llama (RMSE FT Cycle 1 and RMSE FT 2, respectively). The Difference Cycle 1 and Difference Cycle 2 columns indicate the changes in RMSE relative to Optuna for the first and second fine-tuning cycles, respectively. Positive values in the difference columns represent improvements, while negative values indicate an increase in RMSE.
Model Epochs RMSE Optuna RMSE FT 1 Difference
RNN 1 0.7959 0.7880 0.0079
RNN 2 0.7888 0.7703 0.0185
LSTM 1 0.7691 0.7564 0.0127
LSTM 2 0.7468 0.7478 -0.0010
Llama 1 0.9988 0.9990 -0.0002
Llama 2 0.9987 0.9991 -0.0004
@@
Table 3: Comparison of RMSE values for text generation models from the training dataset during fine-tuning cycle 1. The table highlights the differences in RMSE between the baseline results (Optuna) and the fine-tuned results (FT 1). Positive differences indicate an improvement in RMSE after fine-tuning, while negative differences suggest a potential decline in performance.
Model Epochs RMSE Optuna RMSE FT 2 Difference
InceptionV3 1 0.5906 0.8269 -0.2363
InceptionV3 2 0.4323 0.7687 -0.3364
MaxVit 1 0.8349 0.6430 0.1919
MaxVit 2 0.8297 0.5356 0.2941
VisionTransformer 1 0.7768 0.5884 0.1884
VisionTransformer 2 0.7659 0.5253 0.2406
Computer vision models. Table 2 presents a detailed comparison of RMSE values for each model and epoch under three conditions: baseline Optuna, Code Llama after the first fine-tuning cycle (FT Cycle 1), and Code Llama after the second fine-tuning cycle (FT Cycle 2). The final two columns (Difference 1 and Difference 2) show how much RMSE changes relative to the baseline, with positive values indicating that Code Llama yields a lower RMSE than Optuna (i.e., an improvement), and negative values implying the baseline outperforms Code Llama.
A number of architectures exhibit immediate gains in the first fine-tuning cycle. For example, DenseNet, MobileNetV2, RegNet, ShuffleNet, SqueezeNet, SwinTransformer, and VGG consistently show positive differences across their epochs, indicating a clear reduction in RMSE relative to Optuna. ConvNeXt also demonstrates improvements in both epochs, confirming Code Llama’s ability to quickly adapt to a variety of high-performance models. AlexNet shows a notable positive difference in epoch 2, reflecting early and meaningful gains in hyperparameter recommendations. In contrast, some models — such as EfficientNet and GoogLeNet (epoch 1) — register negative differences, implying that for certain configurations, Optuna remains more effective in finding lower-RMSE settings.
The second fine-tuning cycle further improves RMSE in several models. AlexNet continues to perform well, with both epochs showing additional gains over the first cycle. Similarly, GoogLeNet and SqueezeNet exhibit further improvements across both epochs, highlighting the benefits of iterative tuning. In models such as ConvNeXt and MobileNetV3 (epoch 2), the second cycle also yields enhanced performance compared to the first, mitigating earlier shortcomings.
However, not all models benefit equally from iterative LLM-based fine-tuning. For instance, MNASNet shows a drop in performance in the second epoch, indicating a potential limitation in its adaptability to the fine-tuning process. Similarly, ResNet and RegNet show a decline in RMSE improvements in the second cycle, pointing to possible instability in the learned hyperparameter patterns. These observations highlight the importance of carefully assessing trade-offs and potential limitations when applying fine-tuning to specific model types.
@@
When comparing the fine-tuning results to the Optuna baseline, it becomes clear that fine-tuning effectively reduces RMSE for the majority of models and epochs — particularly for computationally demanding architectures, such as ConvNeXt, SwinTransformer, and VGG. These results affirm that fine-tuning is a robust optimization strategy capable of surpassing or complementing traditional hyperparameter search methods like Optuna for many computer vision tasks. At the same time, the variability across models and epochs underscores the need for architecture-specific tuning strategies to maximize the benefits of fine-tuning.
Text generation models.
@@
The analysis of RMSE differences for the first fine-tuning cycle, as shown in Table 3, provides detailed insight into the impact of fine-tuning on text generation models. For the RNN model, fine-tuning results in a small but consistent improvement in RMSE—approximately 0.0185 lower than the Optuna baseline at epoch 2 — indicating stable and beneficial adjustments from the LLM.
The LSTM model also shows improvement, reducing RMSE by about 0.0127 at epoch 1 compared to Optuna, although a minor performance drop is observed at epoch 2. This variation suggests that LSTM responds positively to fine-tuning overall, though tuning depth may affect consistency.
In contrast, the Llama model shows a slight decline in performance across both epochs, with RMSE values marginally worse than Optuna. This may point to the challenges of adapting large, complex architectures like Llama using the current tuning strategy.